Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2))
My lesson doesn't tell me how to solve this question, please help...

Answer by ewatrrr(24785)

 
Hi
Find g(2) first...
g(x) = x2 - 6x - 7.
g(2) = 4-12-7 = -15
f(-15) = -15 + 8 = -7